# Grid Labeling: Crowdsourcing Task-Specific Importance from Visualizations

### v1

@@ Abstract
Knowing where people look in visualizations is key to effective design. Yet, existing research primarily focuses on free-viewing-based saliency models— although visual attention is inherently task-dependent. Collecting task-relevant importance data remains a resource-intensive challenge. To address this, we introduce Grid Labeling – a novel annotation method for collecting task-specific importance data to enhance saliency prediction models. Grid Labeling dynamically segments visualizations into Adaptive Grids, enabling efficient, low-effort annotation while adapting to visualization structure. We conducted a human-subject study comparing Grid Labeling with existing annotation methods, ImportAnnots, and BubbleView across multiple metrics. Results show that Grid Labeling produces the least noisy data and the highest inter-participant agreement with fewer participants while requiring less physical (e.g., clicks/mouse movements) and cognitive effort. {CCSXML} <ccs2012> <concept> <concept_id>10003120.10003145.10003146</concept_id> <concept_desc>Human-centered computing Visualization techniques</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10003120.10003145.10011769</concept_id> <concept_desc>Human-centered computing Empirical studies in visualization</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> </ccs2012>
@@ 1 Introduction
However, existing mouse-tracking-based saliency collection methods rely on free-viewing [NMF20, OAH14, KBB17], as they are designed to encourage users to explore and describe an image. However, because visualizations are often used for analytic tasks [AES05], we posit that an effective model for predicting where people look should consider task relevance when identifying regions of importance.
To address this limitation, we contribute Grid Labelinga toolkit to enhance existing saliency prediction models with task-specific annotation. A key advantage of our tool is that it is more resource-efficient than traditional methods, such as eye-tracking or mouse-tracking [BKO17, OAH14, KBB17, NMF20, WWA24].
Grid Labeling segments visualizations into Adaptive Grids that dynamically adjust based on existing graphical elements, making it easily adapted to various visualization sizes and designs. This approach enables participants to identify critical areas simply by clicking relevant grids, eliminating the need for cumbersome mouse interactions, such as free-form drawing to annotate regions [OAH14] or clicking the same regions multiple times [KBB17]. Moreover, Grid Labeling streamlines data collection, reducing the number of participants required to converge to a stable importance map. In a human-subject experiment, we demonstrate that, compared to the two popular approaches ImportAnnots [OAH14] and BubbleView [KBB17], Grid Labeling produces less noisy data with higher levels of agreement between participant responses. Additionally, participants reported lower perceived effort when using our method. We also explore the key distinction between saliency and importance, contributing to differences in annotation duration.
@@ 2 Related Work
Researchers have been leveraging eye tracking methodologies from human perception research to model how people perceive images [SVNEM15, BMFF15, CPS16]. These models help assess the appearance and salience of visual representations, enabling eye movement tracking to understand the perceptual and cognitive mechanisms of scene perception [IKN98] and object detection [BCJL15]. The existing saliency models perform well in naturalistic scenes ; however, there are unique perception rules and cognitive biases in the artificial world of data visualization  [FPS21, CAFG12, PWV18, KAFB24], and, thus, these models do not accurately predict where people would look in visualizations. Visualization researchers have been building visual saliency models geared to visualizations [MHD17, BRB16]. However, these models rely on handcrafted features, making it difficult to generalize to complex visualizations. Additionally, these models cannot incorporate textual information to generate task-specific saliency maps since the prediction is solely based on visual inputs.
With the advent of deep learning, gaze data were used as the ground truth of saliency models [FCB20, SCHE22, WBB24], leading to higher performance in saliency prediction while enabling task-specific saliency [WWA24]. These models usually need large-scale datasets to learn complex patterns. However, gathering precise gaze data is costly and cumbersome, which limits large-scale data collection efforts. Many researchers, therefore, proposed several proxies for eye gaze. WebGaze [PSL16] uses a webcam for cheap and easy deployment in online studies yet suffers from data quality issues due to low-resolution cameras and uncontrolled calibration. Therefore, mouse-(cursor-)based annotation tools [JHDZ15, KBB17, OAH14] were proposed to improve data quality. Among these methods, BubbleView [KBB17] was the most used tool for capturing visual saliency and importance [BKO17, WWA24]. However, BubbleView is primarily designed for exploring images and gathering information, which differs slightly from the goal of capturing perceived importance. As a result, while BubbleView is well-suited for measuring visual saliency, it may not be the best tool for capturing task-specific importance [NMF20]. Built upon these prior approaches’ limitations, our Grid Labeling aims to collect responses that cover all essential areas of the visualization with minimum noise, leading to more efficient data collection.
@@ 3.2 Adaptive Grid
Step 1: Split Regions. We divide the visualization into three regions: text, edge, and background. We filter out the text area with PaddlePaddle OCR [DLG20], followed by the Canny edge detection algorithm [Can86] to extract graphical elements. The remaining tiles not identified as text or edges are labeled as background.
@@
We solve this optimization problem using Integer Linear Programming (ILP) with OR-Tools’ constraint programming solver [PF19]. By enforcing the exact coverage of each tile, we get patches that cover the visualization while respecting the background contents
@@ 4.1 Participants and Design
We conducted a power analysis based on pilot results. Considering the smallest effect size across all comparison metrics (Cohen’s $f$ = 0.2715 for cognitive load), a target sample of 152 participants would yield 80% power to detect an overall difference between annotation methodologies at an $\alpha$ level of 0.05. We recruited participants from Prolific ($M_{age}$ = 38.5, $\sigma=11.9$, 52 females) and compensated them \$12 per hour. We curated a set of 18 visualizations from ChartQA [MLT22] and CharXiv [WXH24], covering a diverse range of chart types (e.g., bar, stacked bar, pie, line, Choropleth map, heatmap, histogram, scatterplot, and contour plots). Each participant was then shown these selected charts in a random order. In a between-subject set-up, participants were also randomly assigned to annotate them via one of four tools: BubbleView, ImportAnnots, and Static/Adaptive Grid.
@@ 4.2 Procedure
The study was conducted as a between-subject experiment. After consenting to the experiment, participants were instructed on how to use the assigned annotation tool (BubbleView, ImportAnnots, Static Grid, or Adaptive Grid). They first solved an example task with a simple bar chart with one of the following prompts: annotate the important area and describe key points, annotate the area minimally required for you to identify the highest value in the chart (results see Section 4.3). Then, they labeled 18 visualizations using the same tool. For ImportAnnots, participants were instructed as annotate important areas related to answering the question. For BubbleView, they were just asked to answer the question following prior work’s design [WWA24]. For Grid Labeling (Adaptive/Static), they were instructed to annotate “Task-Specific Importance.” In the end, they reported the tool’s usability using NASA-TLX [HS88], completed an assessment of their visualization literacy [PO23], and provided demographic information.
@@ 4.4 Results: Methods Comparison
Coverage Efficiency: Annotated Area per Mouse Travel Distance. We measured the coverage efficiency by dividing the annotated area size by the mouse travel distance during annotation. With ANOVA (F=30.33) and post hoc Tukey test ($p<0.001$), we observed a significant difference between Adaptive Grid and Static Grid having higher coverage efficiency than BubbleView and ImportAnnots.
@@
Convergence Speed: Agreement Across Participants. We examined the number of participants needed to achieve 90% similarity with the aggregated mask using five similarity metrics: Spearman’s Rank Correlation [Spe04], Structural Similarity Index [WBSS04], Dice Coefficient [Dic45], Jaccard Index [Jac12], and Kullback-Leibler (KL) Divergence [KL51]. We measured the convergence of 10 different randomized orders of responses for more generalized results with smoother graphs. As shown in Figure 3, the Adaptive Grid and Static Grid generally converge faster than the other tools across most metrics, while the Adaptive Grid was the best performing in all metrics except Spearman’s R.
@@ 5 Discussion & Future Work
We contribute Grid Labeling, an annotation method for efficiently crowdsourcing task-dependent important areas of visualizations. Grid Labeling outperformed other approaches across all metrics, as shown in Grid Labeling: Crowdsourcing Task-Specific Importance from Visualizations. While ImportAnnots [OAH14] had the lowest click count and BubbleView [KBB17] required the least time, Adaptive Grid achieved the highest inter-participant agreement with the fewest participants across multiple metrics (e.g., SSIM, Dice, Jaccard, KL). Meanwhile, Static Grid demonstrated higher usability, as indicated by NASA-TLX [HS88]. These results highlight the potential of Grid Labeling in training task-specific saliency models, minimizing text overemphasis and enhancing predictive accuracy. Considering the trade-offs, we recommend using an Adaptive Grid for maximizing convergence and a Static Grid to improve usability.
Since the present study did not explore why participants labeled certain grids as important and relied on an inter-participant agreement for quality control, future work could investigate the reasoning behind these selections to provide a more high-level, representational explanation, collecting a large-scale dataset and training a task-specific importance prediction model. Additionally, future work could refine Adaptive Grid generation using vision-based LLMs to enhance annotation usability by semantically filtering less important visualization grids.

### v2

@@ Abstract
Knowing where people look in visualizations is key to effective design. Yet, existing research primarily focuses on free-viewing-based saliency models— although visual attention is inherently task-dependent. Collecting task-relevant importance data remains a resource-intensive challenge. To address this, we introduce Grid Labeling – a novel annotation method for collecting task-specific importance data to enhance saliency prediction models. Grid Labeling dynamically segments visualizations into Adaptive Grids, enabling efficient, low-effort annotation while adapting to visualization structure. We conducted a human-subject study comparing Grid Labeling with existing annotation methods, ImportAnnots, and BubbleView across multiple metrics. Results show that Grid Labeling produces the least noisy data and the highest inter-participant agreement with fewer participants while requiring less physical (e.g., clicks/mouse movements) and cognitive effort. An interactive demo and the accompanying dataset are available at https://github.com/jangsus1/Grid-Labeling. {CCSXML} <ccs2012> <concept> <concept_id>10003120.10003145.10003146</concept_id> <concept_desc>Human-centered computing Visualization techniques</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10003120.10003145.10011769</concept_id> <concept_desc>Human-centered computing Empirical studies in visualization</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> </ccs2012>
@@ 1 Introduction
However, existing mouse-tracking-based saliency collection methods rely on free-viewing [NMF20, OAH14, KBB17], as they are designed to encourage users to explore and describe an image. However, because visualizations are often used for analytic tasks [AES05], we posit that an effective model for predicting where people look should consider task relevance when identifying regions of importance. To address this limitation, we contribute Grid Labeling, a toolkit to enhance existing saliency prediction models with task-specific annotation. A key advantage of our tool is that it is more resource-efficient than traditional methods, such as eye-tracking or mouse-tracking [BKO17, OAH14, KBB17, NMF20, WWA24].
Grid Labeling segments visualizations into Adaptive Grids that dynamically adjust based on existing graphical elements, making it easily adapted to various visualization sizes and designs. This approach enables participants to identify critical areas simply by clicking relevant grids, eliminating the need for cumbersome mouse interactions, such as free-form drawing to annotate regions [OAH14] or clicking the same regions multiple times [KBB17]. Moreover, Grid Labeling streamlines data collection, reducing the number of participants required to converge to a stable importance map. In a human-subject experiment, we demonstrate that, compared to the two popular approaches, ImportAnnots [OAH14] and BubbleView [KBB17], Grid Labeling produces less noisy data with higher levels of agreement between participant responses. Additionally, participants reported lower perceived effort when using our method. We also explore the key distinction between saliency and importance, contributing to differences in annotation duration.
@@ 2 Related Work
Researchers have been leveraging eye tracking methodologies from human perception research to model how people perceive images [KNEM15, CPS16]. These models help assess the appearance and salience of visual representations, enabling eye movement tracking to understand the perceptual and cognitive mechanisms of scene perception [IKN98] and object detection [BCJL15]. The existing saliency models perform well in naturalistic scenes; however, there are unique perception rules and cognitive biases in the artificial world of data visualization [CAFG12], and, thus, these models do not accurately predict where people would look in visualizations. Visualization researchers have been building visual saliency models geared to visualizations [MHD18, BRB16]. However, these models rely on handcrafted features, making it difficult to generalize to complex visualizations. Additionally, these models cannot incorporate textual information to generate task-specific saliency maps since the prediction is solely based on visual inputs.
With deep learning, gaze data became the ground truth for saliency models [SCHE23, WBB24], increasing prediction performance and enabling task-specific saliency [WWA24]. These models require large datasets, but collecting accurate gaze data is expensive and cumbersome. To address this, researchers proposed gaze proxies. For example, WebGaze [PSL16] offers low-cost webcam-based data collection for online studies, though it struggles with quality due to low resolution and poor calibration. Therefore, mouse cursor-based annotation tools [JHDZ15, KBB17, OAH14] were proposed to improve data quality. Among these methods, BubbleView [KBB17] was the most widely used tool for capturing visual saliency and importance [BKO17, WWA24]. However, BubbleView is primarily designed for exploring images and gathering information, which differs slightly from the goal of capturing perceived importance. As a result, while BubbleView is well-suited for measuring visual saliency, it may not be the best tool for capturing task-specific importance [NMF20]. Built upon these prior approaches’ limitations, our Grid Labeling aims to collect responses that cover all essential areas of the visualization with minimum noise, leading to more efficient data collection.
@@ 3.2 Adaptive Grid
Step 1: Split Regions. We divide the visualization into three regions: text, edge, and background. We filter out the text area with PaddlePaddle OCR, followed by the Canny edge detection algorithm to extract graphical elements. The remaining tiles not identified as text or edges are labeled as background.
@@
We solve this optimization problem using Integer Linear Programming (ILP) with OR-Tools’ constraint programming solver. By enforcing the exact coverage of each tile, we get patches that cover the visualization while respecting the background contents.
@@ 4.1 Participants and Design
We conducted a power analysis based on pilot results. Considering the smallest effect size across all comparison metrics (Cohen’s $f$ = 0.2715 for cognitive load), a target sample of 152 participants would yield 80% power to detect an overall difference between annotation methodologies at an $\alpha$ level of 0.05. We recruited participants from Prolific ($M_{age}$ = 38.5, $\sigma=11.9$, 52 females) and compensated them \$12 per hour. We curated a set of 18 visualizations from ChartQA [MLT22] and CharXiv [WXH24], covering a diverse range of chart types (e.g., bar, stacked bar, pie, line, Choropleth map, heatmap, histogram, scatterplot, and contour plots). Each participant was then shown these selected charts in a random order. In a between-subject set-up, participants were randomly assigned to annotate them via one of four tools: BubbleView, ImportAnnots, and Static/Adaptive Grid.
@@ 4.2 Procedure
The study was conducted as a between-subject experiment. After consenting to the experiment, participants were instructed on how to use the assigned annotation tool (BubbleView, ImportAnnots, Static Grid, or Adaptive Grid). They first solved an example task with a simple bar chart with one of the following prompts: annotate the important area and describe key points, annotate the area minimally required for you to identify the highest value in the chart (results see Section 4.3). Then, they labeled 18 visualizations using the same tool. For ImportAnnots, participants were instructed as annotate important areas related to answering the question. For BubbleView, they were just asked to answer the question following the prior work’s design [WWA24]. For Grid Labeling (Adaptive/Static), they were instructed to annotate “Task-Specific Importance.” In the end, they reported the tool’s usability using NASA-TLX [HS88], completed an assessment of their visualization literacy [PO23], and provided demographic information.
@@ 4.4 Results: Methods Comparison
Coverage Efficiency: Annotated Area per Mouse Travel Distance. We measured the coverage efficiency by dividing the annotated area size by the mouse travel distance during annotation. With ANOVA (F=30.33) and post hoc Tukey ($p<0.001$), we observed a significant difference between Adaptive Grid and Static Grid, having higher coverage efficiency than BubbleView and ImportAnnots.
@@
Convergence Speed: Agreement Across Participants. We examined the number of participants needed to achieve 90% similarity with the aggregated mask using five similarity metrics: Spearman’s Rank Correlation [Spe04], Structural Similarity Index [WBSS04], Dice Coefficient [Dic45], Jaccard Index [Jac12], and Kullback-Leibler (KL) Divergence [KL51], often used to measure difference between continuous 2d maps. We measured the convergence of 10 different randomized orders of responses for more generalized results with smoother graphs. As shown in Figure 3, the Adaptive Grid and Static Grid generally converge faster than the other tools across most metrics, while the Adaptive Grid was the best performing in all metrics except Spearman’s R.
@@ 5 Discussion & Future Work
We contribute Grid Labeling, an annotation method for efficiently crowdsourcing task-dependent important areas of visualizations. Grid Labeling outperformed other approaches across all metrics, as shown in Grid Labeling: Crowdsourcing Task-Specific Importance from Visualizations. While ImportAnnots [OAH14] had the lowest click count and BubbleView [KBB17] required the least time, Adaptive Grid achieved the highest inter-participant agreement with the fewest participants across multiple metrics (e.g., SSIM, Dice, Jaccard, KL). Meanwhile, Static Grid demonstrated higher usability, as indicated by NASA-TLX [HS88]. These results highlight the potential of Grid Labeling in training task-specific saliency models, minimizing text overemphasis, and enhancing predictive accuracy. Considering the trade-offs, we recommend using an Adaptive Grid for maximizing convergence and a Static Grid to improve usability.
Since the present study did not explore why participants labeled certain grids as important and relied on an inter-participant agreement for quality control, future work could investigate the reasoning behind these selections to provide a more high-level, representational explanation, collecting a large-scale dataset and training a task-specific importance prediction model. Additionally, future work could refine Adaptive Grid generation using vision-based LLMs to enhance annotation usability by semantically filtering less important visualization grids through an automated pipeline.
Acknowledgements
This work was supported by the National Science Foundation under grants IIS-2237585 and IIS-2311575.